# Rough or crumpled: Strong coupling phases of a generalized Kardar-Parisi-Zhang surface

### v1

@@ Appendix B MCT Calculations
After shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$ above integral reduces to
$\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}(q+\frac{k}{2})_{m}(\frac{k}{2}-q)_{m}(q-\frac{k}{2})_{i}(q-\frac{k}{2})_{j}(q+\frac{k}{2})_{i}\times\frac{{\bf(q+\frac{k}{2})}^{-2\chi-d}}{({\bf q-\frac{k}{2}})^{2}\times\biggl{(}({\bf q+\frac{k}{2}})^{z}+{\bf|q-\frac{k}{2}|}^{z}\biggl{)}}.$ (52)
Next we find ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from numerator, which are relevant to us. Then the above integral $I$ can be split into three parts say $I_{1}$, $I_{2}$ and $I_{3}$. The first part $I_{1}$, after expanding its denominator binomially, is
$I_{1}$ $=-\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{2q^{z}}\times\frac{k_{i}q_{i}}{q^{4}}$
$=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (53)
@@
$I_{2}$ $=\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}\times\frac{k_{i}q_{i}}{q^{2}}$
$=\frac{\lambda\lambda_{1}D}{2\Gamma}(2\chi+d)\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (54)
@@
$I_{3}$ $=\frac{2\lambda\lambda_{1}D}{2\Gamma}k^{2}\int_{{\bf q}}\frac{q^{4}q^{-2\chi-d}}{q^{2}\times 2q^{z}}$
$=\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k_{d}k^{2-\chi}}{\chi}.$ (55)
@@
$-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}\biggl{[}\frac{2-2\chi-2d}{2}\biggl{]}.$ (56)
@@
$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{\chi\times d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (58)
@@
Figure 4(a) gives the one-loop correction at ${\cal O}(\lambda^{2})$ to the correlation function with a symmetry factor of $2$. Evaluating it at zero frequency, we get
@@
$=\frac{\lambda^{2}}{2\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}4D^{2}\Gamma^{2}q^{-2\chi-d+z}\times\frac{|{\bf k-q}|^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(\Omega^{2}+\Gamma^{2}({\bf k-q})^{2z})}$ (59)
@@
After performing the $\Omega$ integral, we find
@@
$=\frac{\lambda^{2}D^{2}}{2\Gamma^{3}}k_{d}\frac{1}{(d+3\chi-2)}k^{-d-\chi-2}.$ (62)
Next, Fig. 4(b) is the one-loop correction at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of $2$. Evaluating it at zero frequency, we get
@@
Using, $\int_{\Omega}\frac{1}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}=\frac{1}{4\Gamma^{3}q^{3z}}$ above integral reduces to,
@@
By using $\chi+z=2$ and identity (40), we get the one-loop contribution to the correlation function at ${\cal O}(\lambda_{1})^{2}$:
@@
Finally, Fig. 4(c) gives the one-loop contribution to zero frequency correlation function at ${\cal O}(\lambda_{1}\lambda)$ with a symmetry factor of $2$. Evaluating, we find
@@
$A=36\gamma^{2}+(6d+30)\gamma+(3d+6)$
$B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)}$

### v3

@@ III MCT Calculations
After shifting ${\bf q}\rightarrow{\bf q}+\frac{\bf k}{2}$, above integral reduces to
$\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}(q+\frac{k}{2})_{m}(\frac{k}{2}-q)_{m}(q-\frac{k}{2})_{i}(q-\frac{k}{2})_{j}(q+\frac{k}{2})_{i}\times\frac{{\bf(q+\frac{k}{2})}^{-2\chi-d}}{{({{\bf q}-\frac{\bf k}{2}})^{2}\biggl{(}|{\bf q}+\frac{\bf k}{2}}|^{z}+|{\bf q}-\frac{\bf k}{2}|^{z}\biggl{)}}.$ (58)
Next we calculate the ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from numerator, which are relevant. Then the above integral $I$ can be split into three parts say $I_{1}$, $I_{2}$ and $I_{3}$. The first part $I_{1}$, after expanding its denominator binomially, is
$I_{1}$ $=-\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{2q^{z}}\,\frac{k_{i}q_{i}}{q^{4}}=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (59)
@@
$I_{2}$ $=\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}\times\frac{k_{i}q_{i}}{q^{2}}=\frac{\lambda\lambda_{1}D}{2\Gamma}(2\chi+d)\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (60)
@@
$I_{3}$ $=\frac{2\lambda\lambda_{1}D}{2\Gamma}k^{2}\int_{{\bf q}}\frac{q^{4}q^{-2\chi-d}}{2q^{2}q^{z}}=\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k_{d}k^{2-\chi}}{\chi}.$ (61)
@@
${I_{1}+I_{2}+I_{3}}=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}\biggl{[}\frac{2-2\chi-2d}{2}\biggl{]}.$ (62)
@@
$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{\chi\,d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (64)
@@
Figure 5(a) gives the one-loop correction at ${\cal O}(\lambda^{2})$ to the correlation function with a symmetry factor of 222. Evaluating it at zero frequency, we get
@@
$=\frac{\lambda^{2}}{2\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}4D^{2}\Gamma^{2}q^{-2\chi-d+z}\times\frac{|{\bf k-q}|^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(\Omega^{2}+\Gamma^{2}|{\bf k-q}|^{2z})}$ (65)
@@
After performing the $\Omega$-integral, we find
@@
Next, Fig. 5(b) is the one-loop correction at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of 222. Evaluating it at zero frequency, we get
@@
Using, $\int_{\Omega}\frac{1}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}=\frac{1}{4\Gamma^{3}q^{3z}}$, above integral reduces to
@@
By using $\chi+z=2$ and identity (46), we get the one-loop contribution to the correlation function at ${\cal O}(\lambda_{1}^{2})$:
@@
Finally, Fig. 5(c) gives the one-loop contribution to zero frequency correlation function at ${\cal O}(\lambda\lambda_{1})$ with a symmetry factor of 222. Evaluating, we find
@@
$A=36\gamma^{2}+(6d+30)\gamma+(3d+6),$
$B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},$